### v2

# Motivic Homotopy Groups of Spheres and Free Summands of Stably Free Modules

Abstract.
Working over an algebraically closed field of characteristic 0, we show that the motivic stable homotopy groups of the sphere spectrum can be determined entirely from the motivic homotopy groups of the $p$-completed sphere spectra and the motivic cohomology of the ground field, except possibly for the 0 and 1-1-stems. Using this, we show that the complex realization maps from the motivic homotopy groups to the classical stable homotopy groups are an isomorphism in a range of bidegrees. We apply this to deduce that complex realization also induces isomorphisms on unstable homotopy groups for Stiefel varieties $V_{r}(\mathbb{A}^{n}_{k})$ in a range of bidegrees. This allows a complete solution of the question of when the projection map $V_{r}(\mathbb{A}^{n}_{k})\to V_{1}(\mathbb{A}^{n}_{k})$ admits a right inverse. Equivalently, this settles the question of when the universal stably-free module of type (n,n1)(n,n-1) admits a free summand of given rank.
Key words and phrases:
Stably free modules, vector bundles, Stiefel varieties, motivic homotopy groups
2020 Mathematics Subject Classification:
13C10, 14F42, 14J60, 57R22
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2021-02603.
1. Introduction
Let $R$ be a ring, assumed unital and commutative, and $P$ an $R$-module. One says that $P$ is stably free of type $(n,r)$ if there is an isomorphism
$P\oplus R^{n-r}\cong R^{n}.$
In [9], it is proved that a stably free module $P$ of type (24m,24m1)(24m,24m-1) admits a free summand of rank $2$ if $R$ contains a field $k$ of characteristic 0, and of rank $3$ if that field has at most one quadratic extension. The method is to convert the problem to motivic homotopy theory: specifically, the injectivity of a realization map on the homotopy groups of Stiefel varieties. This problem is then solved by using recent excellent progress on the calculation of motivic homotopy sheaves of spheres—notably, [22] and [23]. These calculations of the homotopy sheaves are of a general nature, valid over any field.
A great deal more is known about the motivic homotopy groups over an algebraically closed field of characteristic 0, using the motivic Adams and Adams–Novikov spectral sequences: [8, 13, 15, 26, 14] and elsewhere. The apparent obstacle to using these calculations in our work is that these sequences calculate the homotopy groups of $p$-completed objects, rather than of the sphere spectrum itself. In this paper, we explain how to surmount this obstacle using some algebra and the results of [2], and give a large range of bidegrees $(s,w)$ in which the motivic and classical homotopy groups coincide. Using this, and assuming our rings contain $\bar{\mathbb{Q}}$, we can then push the methods of [9] much further, fully solving a problem implicit in [20] of determining which universal examples of stably free modules admit rank-$r$ free summands.
1.1. Outline
In Section 2, we discuss the Ext-completion $\operatorname{Ext}(\mathbb{Q}/\mathbb{Z},A)$ of an abelian group $A$, which may be viewed as a completion at all primes. We also handle the variant where only a subset of the prime numbers is considered. We make no claims to originality about this section, but we need it as a reference for our arguments concerning the homotopy groups of $p$-completed motivic spectra.
Let $k=\bar{k}$ be an algebraically closed field of characteristic 0. The main body of the paper begins with Theorem 3.2, which says that in nonnegative weights $w$ and aside from the $s=0$-stem, the motivic stable homotopy group $\bm{\pi}_{s,w}(\mathbb{1})(k)$ of the sphere spectrum is isomorphic to the product over all primes $p$ of the groups $\bm{\pi}_{s,w}(\mathbb{1}^{\wedge}_{p})(k)$ of the $p$-completed spheres. A weaker version holds for negative weights: the completions may fail to detect the maximal divisible subgroup, which in most cases coincides with the motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
This implies that the extensive calculations of homotopy groups of the $p$-completed sphere amount to computations of the motivic homotopy groups $\bm{\pi}_{s,w}(\mathbb{1})(k)$ of the sphere spectrum when $s\neq 0$ and $w\geq 0$ (in fact, when $w\geq-1$).
An immediate consequence of Theorem 3.2 is that $\bm{\pi}_{s,w}(\mathbb{1})(k)$ is isomorphic to the classical stable homotopy group of the sphere spectrum, $\pi_{s}(\mathbb{S})$, when $s\neq 0$ and $0\leq w\leq\frac{1}{2}(s+1)$—this relies on isomorphisms of the $p$-completions, which are known from [10] and [25].
There are two ingredients in the proof of Theorem 3.2. First we consider the algebraic properties of the $\operatorname{Ext}$-completions $\operatorname{Ext}(\mathbb{Z}/(p^{\infty}),\bm{\pi}_{s,w}(\mathbb{1})(k))$, which agrees with $\bm{\pi}_{s,w}(\mathbb{1}^{\wedge}_{p})(k)$ except when $s=0$. Finiteness of $\prod_{p}\bm{\pi}_{s,w}(\mathbb{1})(k)$ is now sufficient to prove that the completion map belongs in a split short exact sequence
$0\to D_{s,w}\to\bm{\pi}_{s,w}(\mathbb{1})(k)\to\prod_{p}\bm{\pi}_{s,w}(\mathbb{1}^{\wedge}_{p})(k)\to 0$
where $D_{s,w}$ is the subgroup of divisible elements in $\pi_{s,w}(\mathbb{1})(k)$. Except when $s\in\{0,-1\}$, this kernel is observed to be a $\mathbb{Q}$-vector space. We then rely on [2], which draws on [7], for the identification
$\mathbb{Q}\otimes_{\mathbb{Z}}\bm{\pi}_{s,w}(\mathbb{1})(k)\cong\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$
when $k$ is algebraically closed and $s\not\in\{0,-1\}$.
Having proved that the motivic and classical stable homotopy groups of spheres agree, we then exploit the $\mathbb{P}^{1}$-Freudenthal theorem of [3] and an induction argument to prove that complex realization induces isomorphisms on homotopy groups of Stiefel varieties. Here we suppose $k$ is embedded as a subfield of $\mathbb{C}$. If $V_{r}(\mathbb{A}^{n})$ denotes $\operatorname{GL}(n)/\operatorname{GL}(n-r)$ and $W_{r}(\mathbb{C}^{n})$ denotes its complex analogue U(n)/U(nr)U(n)/U(n-r), then the realization map
(1) $\bm{\pi}_{d+e,e}(V_{r}(\mathbb{A}^{n}))(k)\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
is an isomorphism for a set of values of $(d,e)$ depending on $(n,r)$, and an injection for a larger set. The precise statements are Theorems 4.4 and 4.5.
These theorems combine with [9] to imply that right inverses for projection maps $V_{r}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})\to V_{1}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})$ exist if and only if right inverses exist for the corresponding projections $W_{r}(\mathbb{C}^{n})\to W_{1}(\mathbb{C}^{n})$. Following [1], it is known that the latter exists if and only if $b_{r}\mid n$ where $b_{r}$ is the $r$th James number. We deduce Theorem 5.1, saying the same thing for the Stiefel variety.
Finally, we establish Corollary 5.2, which says that if $R$ is a commutative ring containing an algebraic closure of $\mathbb{Q}$, and if $P$ is a stably free $R$-module satisfying $P\oplus R\cong R^{n}$, and if $b_{r}\mid n$, then $P$ admits a decomposition $R^{r}\oplus Q\cong P$. This settles a question left open by [20, Thm. 6.5] in the case where the ground field contains an algebraic closure of $\mathbb{Q}$.
1.2. Remarks and conventions
We use the symbol $\mathbb{1}$ for the motivic sphere spectrum, and $\mathbb{S}$ for the classical sphere spectrum. This usage is borrowed from [23], and is justified by convenience.
When we work with motivic spaces or spectra over a base field $k$, we use a boldface $\bm{\pi}_{\ast}(X)$ for the homotopy sheaves. If $X$ is a motivic object, then $\pi_{\ast}(X)=\bm{\pi}_{\ast}(X)(k)$ is used for the group (or pointed set) of global sections.
There are two useful bigrading conventions for motivic spheres, and consequently for homotopy sheaves and groups. There is the original convention:
$S^{s,w}=S^{s-w}\wedge(\mathbb{A}^{1}_{k}\setminus\{0\})^{\wedge w},$
which we will call the “stem-weight” bigrading, and the convention of [12]:
$S^{c+w\alpha}=S^{c}\wedge(\mathbb{A}^{1}_{k}\setminus\{0\})^{\wedge w},$
which we will call the “coweight-weight” bigrading. Stem-weight bigrading is particularly well suited to motivic Adams and Adams–Novikov spectral sequence calculations over $\mathbb{C}$, and is what is used in the extensive literature on this topic. On the other hand, coweight-weight bigrading is better for use in problems concerning motivic connectivity, for instance, the motivic Freudenthal suspension theorem of [3]. As a consequence, we use both conventions in this paper. In practice, this means that the notation $c+w\alpha$ should be read as $(c+w,w)$. Here is the same homotopy group written in two ways:
$\pi_{c+w\alpha}(X)=\pi_{c+w,w}(X).$
The symbol $\alpha$ does not appear with any other sense in this paper.
Some of our later results, notably Proposition 4.3 and Theorem 4.4, are concerned with maps that are isomorphisms modulo uniquely divisible subgroups of the source. We devote some attention to the homological algebra of such morphisms in Appendix A.
Some of our sources, particularly [8], [10], [14] and other works on the motivic Adams spectral sequence, work over the ground field $\mathbb{C}$. Using [21], for the purposes of calculating $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$, we can safely replace $\mathbb{C}$ by an arbitrary algebraically closed field $k$ of characteristic 0. We ultimately wish to apply our calculations, in Section 5, in the case where $k=\bar{\mathbb{Q}}$, which is the most general available to us. It imposes essentially no extra cost to work throughout over a general algebraically closed field $k$ of characteristic 0.
The calculations of homotopy groups in this paper all hold over an algebraically closed field $k$ of characteristic $\ell>0$, at the possible cost of replacing $\pi_{s,w}(X)$ by $\mathbb{Z}[\ell^{-1}]\otimes_{\mathbb{Z}}\pi_{s,w}(X)$. In fact, the calculations here, and knowledge of the Milnor–Witt $K$-theory, suffice to determine the analogous calculations over even more general fields. We refer the interested reader to [5, Theorem 1.1].
1.3. Acknowledgements
This paper was conceived of following a talk by Zhouli Xu at the International Workshop on Algebraic Topology in Hangzhou, China 2025. We thank the organizers of this conference for staging it. We also thank Zhouli Xu for helpful conversations, and Dan Isaksen and Zhouli Xu both for informative correspondence on the nature of the motivic homotopy groups.
2. Ext-Completions
We make no claims about the originality of this section, the results of which are elementary and, in many cases, implicit in [24]. It is included here for reference and to set notation.
2.1. Definitions
Let $\mathcal{P}$ denote the set of all prime numbers. For the rest of this subsection, fix a subset $I\subseteq\mathcal{P}$, which may be empty. We say that a positive integer $n$ is a product of primes in $I$ if every prime factor $p$ of $n$ lies in $I$. Note that $1$ is a product of primes in $\emptyset$.
Let $n$ be a positive integer and $A$ an abelian group. Consider the morphism $\times n:A\to A$ given by multiplication by $n$. We say that $A$ is $n$-divisible if $\times n$ is an epimorphism, and uniquely $n$-divisible if it is an isomorphism. We say that $A$ is (uniquely) $I$-divisible if it is (uniquely) $n$-divisible for all $n$ that are products of primes in $I$.
If $\times n$ is a monomorphism, then $A$ is $n$-torsion-free, and an abelian group that is $n$-torsion-free for all $n$ divisible by primes in $I$ is said to be $I$-torsion-free. At the other extreme, if $\times n:A\to A$ is the 0-map, we say that $A$ is $n$-torsion. If $A$ is $n$-torsion for some product $n$ of primes in $I$, then $A$ is said to be of $I$-bounded torsion.
The notation above concerns the group as a whole. Given an element $a\in A$, we say that $a$ is $I$-torsion if it is $n$-torsion for some integer $n$ that is a product of primes in $I$; it is $n$-divisible (in $A$) if the equation $nx=a$ can be solved for $x$ in $A$, and uniquely $n$-divisible the equation has a unique solution; $a$ is (uniquely) $I$-divisible if it is (uniquely) $n$-divisible for all integers $n$ that are products of primes in $I$.
In all cases above, if $I=\mathcal{P}$, then $I$ may be omitted from the notation, so we will write “uniquely divisible”, “bounded torsion”, “torsion-free” and so on.
Suppose $F:\mathrm{Ab}^{\text{op}}\to\mathrm{Ab}$ is a contravariant additive functor. From the previous definitions, it is immediate that
  • If $A$ is uniquely $I$-divisible, then $F(A)$ is uniquely $I$-divisible;
  • If $F$ is left exact (e.g., $F(-)=\operatorname{Hom}(-,B)$ for some $B$), and if $A$ is $I$-torsion-free, then $F(A)$ is $I$-torsion-free.
Write $I^{\prime}=\mathcal{P}\setminus I$. We write $\mathbb{Z}[I^{-1}]$ to denote the subring of $\mathbb{Q}$ consisting of numbers that may be written as $a/b$ where $a$ is an integer and $b$ is a product of primes in $I$. Note that $\mathbb{Z}[\emptyset^{-1}]=\mathbb{Z}$ and $\mathbb{Z}[\mathcal{P}^{-1}]=\mathbb{Q}$. We define $\mathbb{Z}/\left(I^{\infty}\right)$ as the quotient group:
$0\to\mathbb{Z}\to\mathbb{Z}[I^{-1}]\to\mathbb{Z}/\left(I^{\infty}\right)\to 0.$
The group $\mathbb{Z}/\left(I^{\infty}\right)$ is torsion, and the torsion order of each element a product of primes in $I$. In the case where $I=\{p\}$ is a singleton, we generally write $\mathbb{Z}[p^{-1}]$ and $\mathbb{Z}/\left(p^{\infty}\right)$. If $I=\mathcal{P}$ is the set of all prime numbers, then we will write $\mathbb{Q}/\mathbb{Z}$ for $\mathbb{Z}/[\mathcal{P}^{\infty}]$.
About these groups, we observe
  • $\mathbb{Z}[I^{-1}]$ is uniquely $I$-divisible.
  • $\mathbb{Z}/\left(I^{\infty}\right)$ is $I$-divisible, and uniquely $I^{\prime}$-divisible: for $I$, this follows from the same fact for $\mathbb{Z}[I^{-1}]$. Suppose $q$ is a prime not in $I$. For any $x\in\mathbb{Z}/\left(I^{\infty}\right)$, we can find some product $n$ of primes in $I$ such that $nx=0$. Then $q$ divides $n^{q-1}-1$, so that $\frac{1-n^{q-1}}{q}$ is an integer. If we multiply $\frac{1-n^{q-1}}{q}x$ by $q$, we obtain $x$, so that $\times q$ is indeed surjective. Injectivity of $\times q$ follows from this: for any $y\in\mathbb{Z}[I^{-1}]$, the product $qy$ is an integer if and only if $y$ is.
We deduce that, for any abelian group $A$:
  • $\operatorname{Hom}(\mathbb{Z}[I^{-1}],A)$ and $\operatorname{Ext}(\mathbb{Z}[I^{-1}],A)$ are uniquely $I$-divisible.
  • $\operatorname{Hom}(\mathbb{Z}/\left(I^{\infty}\right),A)$ and $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ are uniquely $I^{\prime}$-divisible.
  • $\operatorname{Hom}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is torsion-free.
Definition 2.1 ([6]).
The $\operatorname{Ext}$-completion of an abelian group $A$ at the set of primes $I$ is the group $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$.
We have observed above that the $\operatorname{Ext}$-completion consists of nontorsion or elements whose torsion order is a product of primes in $I$. From the short exact sequence
(2) $0\to\mathbb{Z}\to\mathbb{Z}[I^{-1}]\to\mathbb{Z}/\left(I^{\infty}\right)\to 0$
we deduce that there is a natural transformation (a connecting homomorphism)
$A=\operatorname{Hom}(\mathbb{Z},A)\to\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A),$
and we will refer to this as the “$I$-completion map” in the sequel, or the “$p$-completion map” when $I=\{p\}$. As usual, if $I=\mathcal{P}$, the symbol “II” may be omitted.
If $I_{0}\subseteq I$ is a subset of $I$, then there is a canonical homomorphism $\mathbb{Z}/(I_{0}^{\infty})\to\mathbb{Z}/(I^{\infty})$. In particular, for any set of prime numbers, we obtain a canonical homomorphism
(3) $\bigoplus_{p\in I}\mathbb{Z}/\left(p^{\infty}\right)\to\mathbb{Z}/\left(I^{\infty}\right).$
Proposition 2.2.
The canonical homomorphism of (3) is an isomorphism.
Proof.
Surjectivity of the homomorphism follows from surjectivity of $\bigoplus_{p\in I}\mathbb{Z}[p^{-1}]\to\mathbb{Z}[I^{-1}]$, where it follows from the existence of partial fraction decompositions for rational numbers.
For injectivity, suppose that some sequence
$\left(\frac{a_{1}}{p_{1}^{s_{1}}},\frac{a_{2}}{p_{2}^{s_{2}}},\dots,\frac{a_{r}}{p_{r}^{s_{r}}}\right)\in\bigoplus_{p\in I}\frac{\mathbb{Z}}{(p^{\infty})}$
maps to 0 under the canonical homomorphism, where $0\leq a_{i}<p_{i}^{s_{i}}$ for all $i$. That is, suppose
$\sum_{i=1}^{r}\frac{a_{i}}{p_{i}^{s_{i}}}\text{\qquad is an integer.}$
By clearing denominators and considering the power of each prime dividing the numerator, we deduce that each $a_{i}$ must be 0. This establishes injectivity. ∎
We have defined the $I$-completion for our own convenience. The next corollary shows that it is sufficient only to define $p$-completions. The proof is that $\operatorname{Ext}$ converts sums in the first variable into products.
Corollary 2.3.
Let $A$ be an abelian group and $I$ a set of prime numbers. Then the canonical isomorphism $\bigoplus_{p\in I}\mathbb{Z}/\left(p^{\infty}\right)\to\mathbb{Z}/\left(I^{\infty}\right)$ induces an isomorphism
$\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)\to\prod_{p\in I}\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),A).$
2.2. Surjectivity of completion
Proposition 2.4.
Let $A$ be an abelian group and $I$ a set of prime numbers. Suppose $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is a torsion group. Then the completion $A\to\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is surjective.
Proof.
From the exact sequence (2), we derive a long exact sequence
(4) $\cdots\to\operatorname{Hom}(\mathbb{Z}[I^{-1}],A)\to A\to\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)\overset{j}{\to}\operatorname{Ext}(\mathbb{Z}[I^{-1}],A)\to\operatorname{Ext}(\mathbb{Z},A)=0.$
It suffices to show that the map $j$ is 0. The target is uniquely $I$-divisible, so that $\times n^{-1}$ is a well defined automorphism of $\operatorname{Ext}(\mathbb{Z}[I^{-1}],A)$ whenever $n$ is a product of primes in $I$. By hypothesis the source consists of torsion elements, and orders of these elements are products of primes in $I$. For any such $a\in\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$, we can find some $n$ for which $na=0$. Then $j(na)=0$, but where $nj(a)=0$ implies $j(a)=0$. ∎
In the case where $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is a group of bounded torsion, for instance when it is finite, we can say more. The argument involves a step that we set aside as a lemma for later use.
Lemma 2.5.
Let $I$ be a set of prime numbers and $n$ a product of primes in $I$. Let
$0\to A\to B\to C\to 0$
be a short exact sequence of abelian groups in which $A$ is uniquely $I$-divisible and $C$ is $n$-torsion. Then the short exact sequence is split.
Proof.
Consider $\operatorname{Ext}(C,A)$, on which $\times n$ acts as an isomorphism, since $A$ is uniquely $I$-divisible, and as 0, since $C$ is $n$-torsion. It must be the case that $\operatorname{Ext}(C,A)=0$. The Yoneda interpretation of $\operatorname{Ext}$ implies the sequence splits. ∎
Proposition 2.6.
Let $A$ be an abelian group and $I$ a set of prime numbers. Suppose $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is a group of $I$-bounded torsion and $\operatorname{Hom}(\mathbb{Z}/\left(I^{\infty}\right),A)=0$. There is a short exact sequence
[…]
and a splitting map $s$. In this diagram, $j$ is the completion map, the image of $g$ is exactly the subgroup of $I$-divisible elements in $A$, and the image of $s$ is the $I$-torsion subgroup of $A$.
Proof.
Since $\operatorname{Hom}(\mathbb{Z}/\left(I^{\infty}\right),A)=0$, the long exact sequence (4) contains the short exact sequence (5), using Proposition 2.4 for exactness on the right. The group $\operatorname{Hom}(\mathbb{Z}[I^{-1}],A)$ is uniquely $I$-divisible, whereas $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is of $I$-bounded torsion by hypothesis. Lemma 2.5 tells us that a splitting map $s$ can be constructed.
Let $n$ be a product of primes in $I$ with the property that $\times n$ annihilates $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)\cong\operatorname{Im}(s)$, and suppose $a$ is $I$-divisible in $A$. We can find some $b$ for which $nb=a$, and we may decompose $b$ uniquely as $b=g(x)+s(y)$. Then $a=nb=g(nx)+s(ny)=g(nx)\in\operatorname{Im}(g)$. Therefore $\operatorname{Im}(g)$ contains all the $I$-divisible elements of $A$. It is itself $I$-divisible, so it is precisely the subgroup of $I$-divisible elements in $A$.
Finally, since $g$ is injective, and $\operatorname{Hom}(\mathbb{Z}[I^{-1}],A)$ is $I$-torsionfree, we see that every $I$-torsion element of $A$ must lie in $\operatorname{Im}(s)$. ∎
3. The sphere spectrum over $k$
We continue to work over an algebraically closed field $k$ of characteristic 0. The previous results on completions apply well to the motivic sphere spectrum, $\mathbb{1}$. If $p$ is a prime number and $X$ is a motivic spectrum, then [21, §3] constructs the $p$-completion $X\to X^{\wedge}_{p}$. There are exact sequences
(6) $0\to\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{r,s}(X))\overset{i}{\to}\pi_{r,s}(X^{\wedge}_{p})\overset{j}{\to}\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{r-1,s}(X))\to 0.$
Relying on [21], we know that the groups $\pi_{r,s}(X^{\wedge}_{p})$ are independent of the particular choice of $k$. These groups are intensively studied, and the following result is well known to the experts in the field. We learned the proof from Dan Isaksen and Zhouli Xu.
Proposition 3.1.
Let $s$ be an integer different from 0 and let $w$ be an integer. As $p$ ranges over the prime numbers, the groups $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ are finite, and almost all of them are 0.
Proof.
For any given weight $w$, there is an Adams–Novikov spectral sequence converging to $\pi_{*,w}(\mathbb{1}^{\wedge}_{p})$: this is constructed for $p=2$ in [12] and in [25] for odd primes (see also [26]). The $E_{2}$-page of this spectral sequence is isomorphic to a bigraded subgroup of the $E_{2}$-page of the classical Adams–Novikov spectral sequence. This follows from [12, Equation (36)] in the case of $p=2$, and from [25, Prop. 2.4.4(2)]. The classical vanishing results [19, Cor. 2.1, Cor 3.1] imply that for each stem $s$, there are only finitely many primes $p$ for which these groups are not 0, and for each such prime $p$, there are only finitely many associated nonzero groups on the $E_{2}$-page and these are themselves finite abelian groups. Since the $E_{\infty}$-page of the motivic Adams–Novikov spectral sequence is a subquotient of the $E_{2}$-page, it follows that $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ has a finite filtration whose associated graded groups are finite. In particular, $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ is a finite abelian group. ∎
Theorem 3.2.
Let $s,w$ be integers satisfying $s\neq 0$. Then the comparison map
$\pi_{s,w}(\mathbb{1})\to\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$
is split surjective, and the kernel is the subgroup of divisible elements. Additionally, if $s\neq-1$, then the kernel is isomorphic to the motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
Proof.
We remark that the motivic cohomology groups $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ are uniquely divisible except when $w\leq 0$ and s=0,1s=0,-1. This is a consequence of the long exact sequence
$\cdots\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))\overset{\times n}{\to}\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}/(n)(-w))\to\mathrm{H}^{-s+1}(\operatorname{Spec}k;\mathbb{Z}(-w))\to\cdots$
and the norm-residue isomorphism theorem ([11, Thm. 1.8]) for the algebraically closed field $k$. As a result, rationalization induces an isomorphism
$\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w)),\text{\quad provided $s\neq 0,-1$.}$
We use this to replace the rational cohomology by the integral at the end of the proof.
When $s\neq 0$, the group $\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ is finite by Proposition 3.1. In particular, it is $n$-torsion for some positive integer $n$.
For a fixed $p$, in the exact sequence (6), we remark that $\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s-1,w}(\mathbb{1}))$ is a torsion-free quotient of a torsion group, so it vanishes, and therefore $\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s,w}(\mathbb{1}))\cong\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$. From this, Corollary 2.3 tells us that the comparison
$\operatorname{Ext}(\mathbb{Q}/\mathbb{Z},\pi_{s,w}(\mathbb{1}))\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$
is an isomorphism.
Since this group is finite, Proposition 2.4 applies, saying that the completion map is surjective. When $s\neq-1$, we have also seen that $\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s+1,w}(\mathbb{1}))=0$, so that Proposition 2.6 applies, telling us the completion is split surjective and the kernel is the subgroup of divisible elements.
In the case of s=1s=-1, a different argument is required. In this case, the Adams and Adams–Novikov spectral sequences imply (see [10] and [25, Prop. 2.4.8]) that $\pi_{-1,w}(\mathbb{1}^{\wedge}_{p})=0$ for all primes $p$. Therefore the completion map is split surjective for trivial reasons. We also see that $\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{-1,w}(\mathbb{1}))=0$, by short exact sequence (6). We may, however, have a nontrivial presentation arising from exact sequence (4):
$\operatorname{Hom}(\mathbb{Q}/\mathbb{Z},\pi_{-1,w}(\mathbb{1}))\to\operatorname{Hom}(\mathbb{Q},\pi_{-1,w}(\mathbb{1}))\to\pi_{-1,w}(\mathbb{1})\to 0.$
In this case we may conclude that $\pi_{-1,w}(\mathbb{1})$ is a quotient of a uniquely divisible group, and so is divisible.
For the rest of this proof, we assume $s\neq-1$. The comparison map belongs in a (split) short exact sequence:
$0\to K_{s,w}\to\pi_{s,w}(\mathbb{1})\to\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})\to 0$
where the kernel $K_{s,w}$ is uniquely divisible, i.e., a $\mathbb{Q}$-vector space, and the cokernel is finite. Therefore, applying $\mathbb{Q}\otimes_{\mathbb{Z}}$- to the short exact sequence yields an isomorphism
(7) $K_{s,w}\cong\mathbb{Q}\otimes_{\mathbb{Z}}\pi_{s,w}(\mathbb{1}).$
Except in the case where $s=w$, an isomorphism $\mathbb{Q}\otimes_{\mathbb{Z}}\pi_{s,w}(\mathbb{1})\cong\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))$ is contained in [2, Thm. 3.7].
We finally consider the case of $s=w$. Here we know the groups $\pi_{s,s}(\mathbb{1})\cong{{\mathbf{K}}}^{\mathsf{MW}}_{-s}(k)$, from [18, Thm. 1.23, Cor. 1.25]. There is a presentation of ${{\mathbf{K}}}^{\mathsf{MW}}_{-s}(k)$ as a product of the Milnor $K$-theory group ${{\mathbf{K}}}^{\mathsf{M}}_{-s}(k)$ with a power $I^{-s}(k)$ of the fundamental ideal of the Witt ring of the field, fibred over $\mathbf{k}^{\mathsf{M}}_{-s}(k)$ the mod-$2$ reduction of the Milnor $K$-theory: see [18, Lem. 3.10]. Over an algebraically closed field, the fundamental ideal vanishes, as does the mod-$2$ Milnor $K$-theory. We are left with an isomorphism ${{\mathbf{K}}}^{\mathsf{MW}}_{-s}(k)\cong{{\mathbf{K}}}^{\mathsf{M}}_{-s}(k)$. The isomorphism $\mathbb{Q}\otimes_{\mathbb{Z}}{{\mathbf{K}}}^{\mathsf{M}}_{-s}(k)\cong\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-s))$ follows from [17, Ch. 5]. ∎
Corollary 3.3.
Let $s,w$ be integers satisfying $s\neq 0$, $w\geq-1$ and $(s,w)\neq(-1,-1)$. The comparison map
$\pi_{s,w}(\mathbb{1})\to\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$
is an isomorphism.
Proof.
First, suppose $s\neq-1$ and $w\geq-1$. Then Theorem 3.2 tells us the map is surjective and the kernel is a motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$. This group vanishes by [17, Ch. 4].
Now suppose s=1s=-1 and $w\geq 0$. Here $\pi_{-1,w}(\mathbb{1})=\pi_{-1-w+w\alpha}(\mathbb{1})=0$ by Morel’s calculations [18, Cor. 6.43], so that surjectivity of the comparison map implies isomorphism. ∎
Recall that we write $\mathbb{S}$ for the sphere spectrum in classical topology. For any $s,w$, there are complex realization maps
$\pi_{s,w}(\mathbb{1})\to\pi_{s}(\mathbb{S})$
and similarly for the $p$-completions.
Corollary 3.4.
Suppose $k$ is a subfield of $\mathbb{C}$. Let $s,w$ be integers satisfying $s\neq 0$ and $w\leq\frac{1}{2}s+1$. Then the complex realization
$\pi_{s,w}(\mathbb{1})\to\pi_{s}(\mathbb{S})$
is split surjective and the kernel is the subgroup of divisible elements. If additionally $s\neq-1$, the kernel is isomorphic to the motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
Proof.
For all $s$ and $w$, the following diagram commutes, disregarding the $\cong$” symbols,
$\pi_{s,w}(\mathbb{1})$$\pi_{s}(\mathbb{S})$$\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$$\prod_{p\in\mathcal{P}}\pi_{s}(\mathbb{S}^{\wedge}_{p}),$$\scriptstyle{\cong}\scriptstyle{\cong}$
where the horizontal arrows are given by complex realization and the vertical arrows are completions. When $s\neq 0$, connectivity of the sphere spectrum and Serre’s thesis imply that $\pi_{s}(\mathbb{S})$ is finite so that the right vertical arrow is an isomorphism. When $w\leq\frac{1}{2}s+1$, [10] (for the prime $2$) and [25] (for odd primes) tell us that the lower horizontal map is an isomorphism.
Therefore, when $s\neq 0$ and $w\leq\frac{1}{2}s+1$, whatever properties enjoyed by the completion map $\pi_{s,w}(\mathbb{1})\to\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ are also enjoyed by the complex realization map. The result now follows from Theorem 3.2. ∎
Remark 3.5.
Note that Corollary 3.4 assures us that the complex realization is an isomorphism in many cases, since $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))=0$ when $w\geq-1$ and $s\neq 0,-1$.
The Beilinson–Soulé conjecture holds that $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))\cong 0$ when $s\geq 1$, independently of $w$. This conjecture is known to hold over number fields, and by [17, Lem. 3.8], it therefore holds for $k=\bar{\mathbb{Q}}$. Therefore we can say that the realization $\pi_{s,w}(\mathbb{1}_{\bar{\mathbb{Q}}})\to\pi_{s}(\mathbb{S})$ is an isomorphism whenever $w\leq\frac{1}{2}s+1$ and $s\geq 1$.
Remark 3.6.
The reason for our preference of the integral cohomology $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ to the rational cohomology $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))$ in Theorem 3.2 is the following observation. Corollary 3.4 holds in full even when (s,w)=(1,1)(s,w)=(-1,-1) and (s,w)=(1,2)(s,w)=(-1,-2) when integer coefficients are used. In both cases, the classical homotopy group is $\pi_{-1}(\mathbb{S})=0$, and the calculations due to [18, Cor. 6.43] and [23, Table 1] give us
$\pi_{-1,-1}(\mathbb{1})\cong{{\mathbf{K}}}^{\mathsf{MW}}_{1}(k)\cong{{\mathbf{K}}}^{\mathsf{M}}_{1}(k)\cong\mathrm{H}^{1}(\operatorname{Spec}k;\mathbb{Z}(1))$
and
$\pi_{-1,-2}(\mathbb{1})\cong\mathrm{H}^{1}(\operatorname{Spec}k;\mathbb{Z}(2)).$
Note that in [23, Table 1], all the other terms vanish over an algebraically closed field of characteristic 0, and the spectral sequence concerned collapses at the $\mathrm{E}_{2}$-stage.
We conjecture that the kernel of the realization map is isomorphic to $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ for an algebraically closed field $k$, provided $s\neq 0$.
Remark 3.7.
If $\bar{k}$ is not a subfield of $\mathbb{C}$, but is an algebraically closed field of characteristic 0, a result similar to Corollary 3.4 holds, by virtue of [21]. The realization map may be replaced by a zigzag $\pi_{s,w}(\mathbb{1})\leftarrow\pi_{s,w}(\mathbb{1}_{\bar{\mathbb{Q}}})\to\pi_{s}(\mathbb{S})$, where $\mathbb{1}_{\bar{\mathbb{Q}}}$ denotes the motivic sphere spectrum over $\bar{\mathbb{Q}}$. This zigzag consists of isomorphisms when $0\leq w\leq\frac{1}{2}s+1$. When $w<0$ and $s\not\in\{-1,0\}$, we can say that $\pi_{s,w}(\mathbb{1})$ admits a decomposition as a direct sum of a finite group isomorphic to $\pi_{s}(\mathbb{S})$ and a uniquely divisible group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
Remark 3.8.
We write the bounds of Corollary 3.4 in coweight-weight terms for future reference. Let $c,w$ be integers satisfying $c+w\neq 0$ and $w\leq c+2$. Then the realization map
$\pi_{c+w\alpha}(\mathbb{1})\to\pi_{c+w}(\mathbb{S})$
is split surjective and the kernel is the subgroup of divisible elements. Except when c+w=1c+w=-1, this kernel is isomorphic to the motivic cohomology group $\mathrm{H}^{-c-w}(\operatorname{Spec}k;\mathbb{Z}(-w))$. In particular, the realization map is an isomorphism when $w\geq-1$, and if the Beilinson–Soulé vanishing conjecture holds for $k$, it is an isomorphism when $c+w>0$.
4. Unstable homotopy groups of spheres and Stiefel varieties
We use the results of [3] to extend Corollary 3.4 to $\pi_{s,w}(S^{a,b})$, the unstable motivic homotopy groups of spheres. For the sake of simplicity of exposition, we assume in this section that the ground field $k=\bar{k}$ is embedded in $\mathbb{C}$.
Recall that there are adjoint functors $\Sigma_{\mathbb{P}^{1}}^{\infty}\dashv\Omega^{\infty}_{\mathbb{P}^{1}}$ between the categories of pointed motivic spaces and motivic $\mathbb{P}^{1}$-spectra. One writes $Q$ for $\Omega^{\infty}_{\mathbb{P}^{1}}\Sigma^{\infty}_{\mathbb{P}^{1}}$. The unstable homotopy sheaves of $QX$ are isomorphic to the stable homotopy sheaves of $X$. There is a unit natural transformation $\mathrm{id}\to Q$, which we call the stabilization.
Proposition 4.1 (Asok–Bachmann–Hopkins, [3]).
Let $a,b$ be nonnegative integers and $s,w$ be integers. Suppose the following inequalities hold:
  1. (1)
    $b\geq 2$ and $a-b\geq 2$;
  2. (2)
    $s-w\leq\min\{a-b-2,b-2\}$;
then the stabilization $S^{a,b}\to QS^{a,b}$ induces an isomorphism of homotopy sheaves
$\bm{\pi}_{a+s,b+w}(S^{a,b})\to\bm{\pi}_{s,w}(\mathbb{1}).$
Proof.
In the notation of [3], $S^{a,b}\in O(S^{a,b})$. Therefore by [3, Theorem 6.3.4], the fibre $\mathcal{F}$ of the stabilization $S^{a,b}\to QS^{a,b}=\Omega^{\infty}_{\operatorname{P}^{1}}\Sigma^{\infty}_{\operatorname{P}^{1}}S^{a,b}$ lies in $O(S^{p,2b})$ where $p$ is the minimum of 2a12a-1 and a+2b1a+2b-1. Then by [3, Lemma 3.1.19], the motivic space $\mathcal{F}$ is p2b1p-2b-1-$\mathbb{A}^{1}$-connected. Recall that $\mathbb{A}^{1}$-connectivity is measured by coweight, see [18, p. 165]. That is to say
(8) $\bm{\pi}_{a+s,b+w}(S^{a,b})\to\bm{\pi}_{s,w}(\mathbb{1})$
is an isomorphism provided
$a-b+s-w\leq\min\{2a-1,a+2b-1\}-2b-1=\min\{2(a-b)-2,a-2\}$
or in other words
$s-w\leq\min\{a-b-2,b-2\},$
which is what we wanted ∎
Remark 4.2.
The result is slightly easier to read when we use coweight-weight grading rather than stem-weight grading. In those terms, it reads as follows: Let $x,y,c,w$ be integers. Suppose the following inequalities hold:
  1. (1)
    $x\geq 2$ and $y\geq 2$;
  2. (2)
    $c\leq\min\{x-2,y-2\}$.
Then the stabilization map
$\bm{\pi}_{x+c+(y+w)\alpha}(S^{x+y\alpha})\to\bm{\pi}_{c+w\alpha}(\mathbb{1})$
is an isomorphism.
The next result is little more than a combination of Corollary 3.3 with Proposition 4.1, but we set it aside as a proposition for later reference.
Proposition 4.3.
Let $d,e,x,y$ be nonnegative integers satisfying:
  1. (1)
    $x\geq 2$ and $y\geq 2$;
  2. (2)
    $d\leq\min\{2x-2,x+y-2\}$;
  3. (3)
    $e-y\leq d-x+2$;
  4. (4)
    $d+e\neq x+y$.
Then the unstable homotopy group $\pi_{d+e}(S^{x+y})$ lies in the stable range and the realization map
$\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y})\cong\pi_{d+e-(x+y)}(\mathbb{S})$
is split surjective and the kernel is the subgroup of arbitrarily divisible elements. Additionally
  • if $d+e\neq x+y-1$, then the kernel is a uniquely divisible group.
  • if $y-1\leq e$ the realization map is an isomorphism. This may be promoted to $\min\{y-1,y+x-d+1\}\leq e$ if the Beilinson–Soulé conjecture holds for $k$.
Proof.
We consider a commutative square in which horizontal arrows are stabilization maps and vertical arrows are $\mathbb{C}$-realization:
$\pi_{d+e\alpha}(S^{x+y\alpha})$$\pi_{(d-x)+(e-y)\alpha}(\mathbb{1})$$\pi_{d+e}(S^{x+y})$$\pi_{d+e-x-y}(\mathbb{S}).$$\scriptstyle{\cong}\scriptstyle{\cong}$
Inequality 2 implies a fortiori that
$d\leq\frac{(2x-2)+(x+y-2)}{2}=\frac{3x+y-4}{2}.$
Combine this with
$d+e\leq d+d-x+y+2=2d-x+y+2$
to deduce
$d+e\leq 2\frac{3x+y-4}{2}-x+y+2=2x+2y-2.$
Therefore the classical Freudenthal suspension theorem applies to say that the lower arrow in the diagram is an isomorphism.
The top arrow is an isomorphism by Proposition 4.1, using inequalities 1 and 2.
The right arrow is split surjective and the kernel is the subgroup of divisible elements, by Corollary 3.4 using inequalities 3 and 4. The two additional claims follow from Corollary 3.4 and Remark 3.5. ∎
To illustrate the result, we provide Figure 1. The figure is supposed to explain what Proposition 4.3 tells us about the complex realization map
$\pi_{d+e,e}(S^{x+y,y})=\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y}).$
We have fixed values for $x,y$ (subject to the condition that $x,y\geq 2$), and indicate regions of interest in the $d,e$-plane.
To the right of the vertical line $d=\min\{2x-2,x+y-2\}$, the motivic Freudenthal suspension theorem does not apply, so our methods break down. Above the diagonal line ey=dx+2e-y=d-x+2, the comparison of the relevant motivic and classical stable homotopy groups of spheres is not understood. When $d+e=x+y$, the stable motivic homotopy group is in the 0-stem, so our results on $\operatorname{Ext}$-completion do not apply. To a lesser extent, this affects the line d+e=x+y1d+e=x+y-1 corresponding to the 1-1-stem, where our understanding of the kernel is worse than elsewhere. To the inequalities of Proposition 4.3 another can be added: when $d<x$, the source group $\pi_{d+e\alpha}(S^{x+y\alpha})$ (corresponding to $\pi_{d-x+(e-y)\alpha}(\mathbb{1})$ of negative coweight) vanishes by a connectivity result of [18].
In Figure 1, we identify four regions in which the comparison map $\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y})$ takes on certain characteristics. To the left of the line $d=x$, the source of the map is 0, and we will call this the “0-region”. Second, marked by vertical hatching, there is a closed pentagonal region labelled with the symbol $\cong$”. This is a region where the map is known to be an isomorphism. Third, there is a triangular region marked with crosshatching and labelled “B–S”. Here the map is an isomorphism contingent on the Beilinson–Soulé vanishing conjecture. This region contains its rightmost edge but not the rest of the boundary. Fourth, marked by horizontal hatching, there is a vertical strip with slanted top, containing its boundary, and labelled with “div”. Here the target of comparison map vanishes, and the group $\pi_{d+e\alpha}(S^{x+y\alpha})$ is known to be divisible, and except possibly along the 1-1-stem line, is isomorphic to $\mathrm{H}^{x+y-d-e}(\operatorname{Spec}k;\mathbb{Z}(y-e))$.
$e$ $d$ $d=\min\{2x-2,x+y-2\}$ey=dx+2e-y=d-x+20-stem,  d+e=x+yd+e=x+y1-1-stem, d+e=x+y1d+e=x+y-1weight 1-1,   e=y1e=y-1coweight 0,  $d=x\congdivB$–S
Figure 1. Behaviour of the comparison $\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y})$ for fixed $x,y$. Points $(d,e)$ in which $d<x$ are not marked, since $\pi_{d+e\alpha}(S^{x+y\alpha})$ vanishes for dimensional reasons.
The case of the spheres is a base case for an induction that allows us to say something about many Stiefel varieties.
Theorem 4.4.
Let $d,e$ be integers and $r,n$ be positive integers. Suppose all the following hold
  1. (1)
    $r\leq n-2$;
  2. (2)
    $d\leq 2n-2r-3$;
  3. (3)
    $e\leq d+4-r$;
  4. (4)
    $2n\leq e+d$;
The realization map
$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
is split surjective, the kernel is the maximal divisible subgroup of the domain, which is uniquely divisible, and the target is a torsion group.
If in addition $n-1\leq e$, then the realization map is an isomorphism.
Proof.
Let us disregard the last paragraph for the moment. The proof is by induction on $r$. The case of $r=1$ follows from Proposition 4.3 with x=n1x=n-1 and $y=n$, since $V_{1}(\mathbb{A}^{n})\simeq S^{n-1+n\alpha}$. The inequalities on $e,d$ imply that $\pi_{d+e}(W_{1}(\mathbb{C}^{n}))=\pi_{d+e}(S^{2n-1})$ is a stable homotopy group of a sphere and inequality 4 implies that it is torsion.
For the induction step, suppose $r\geq 2$ but $r\leq n-2$. We consider the commutative diagram whose rows are exact sequences
(9) $\pi_{d+1+e\alpha}(V_{1}(\mathbb{A}^{n}))$$\pi_{d+e\alpha}(V_{r-1}(\mathbb{A}^{n-1}))$$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))$$\pi_{d+e\alpha}(V_{1}(\mathbb{A}^{n}))$$\pi_{d-1+e\alpha}(V_{r-1}(\mathbb{A}^{n-1}))$$\pi_{d+e+1}(W_{1}(\mathbb{C}^{n}))$$\pi_{d+e}(W_{r-1}(\mathbb{C}^{n-1}))$$\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$$\pi_{d+e}(W_{1}(\mathbb{C}^{n}))$$\pi_{d+e-1}(W_{r-1}(\mathbb{C}^{n-1})).$$\scriptstyle{f_{1}}$$\scriptstyle{f_{2}}$$\scriptstyle{f_{3}}$$\scriptstyle{f_{4}}$$\scriptstyle{f_{5}}$
The commutativity of this diagram is proved in [9, Prop. 5.1]. Map $f_{1}$ is surjective with uniquely divisible kernel by Proposition 4.3: the required inequalities may be easily verified with x=n1x=n-1, $y=n$ and $d+1$ playing the part of $d$. Note that inequalities 14 imply the same inequalities with r1r-1, n1n-1 in place of $r$, $n$. Therefore, the induction hypothesis implies that the map $f_{2}$ is surjective with uniquely divisible kernel. Map $f_{4}$ is surjective with uniquely divisible kernel by Proposition 4.3 again. And map $f_{5}$ is a surjection with uniquely divisible kernel by the induction hypothesis again: inequalities 14 imply the same inequalities with r1r-1, n1n-1 and d1d-1 in place of $r,n$ and $d$. Additionally, the codomains of $f_{2}$ and $f_{4}$ are torsion abelian groups, by the induction hypothesis. It follows that the codomain of $f_{3}$ is also torsion.
It follows from the five-lemma, Proposition A.2, that $f_{3}$ is surjective with uniquely divisible kernel. Lemma 2.5 now assures us that $f_{3}$ is split.
Now we return to the question of isomorphism. If additionally $n-1\leq e$, then all instances of “surjective” may be replaced above with “isomorphic”, by reference to Proposition 4.3, and the usual five-lemma may be used in place of Proposition A.2. ∎
Theorem 4.5.
Let $d,e$ be integers and $r,n$ be positive integers. Suppose all the following hold
  1. (1)
    $r\leq n-2$;
  2. (2)
    $d\leq 2n-2r-3$;
  3. (3)
    $e\leq d+3$;
  4. (4)
    $2n\leq e+d$;
  5. (5)
    $n-1\leq e$.
The realization map
$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
is injective.
Proof.
The proof is by induction on $r$, and follows the proof of Theorem 4.4 extremely closely. In the case of $r=1$, the inequalities above coincide with the inequalities of the stronger form of Theorem 4.4.
For the induction step, suppose $r\geq 2$ but $r\leq n-2$. We consider the commutative diagram whose rows are exact sequences
$\pi_{d+1+e\alpha}(V_{1}(\mathbb{A}^{n}))$$\pi_{d+e\alpha}(V_{r-1}(\mathbb{A}^{n-1}))$$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))$$\pi_{d+e\alpha}(V_{1}(\mathbb{A}^{n}))$$\pi_{d+1+e}(W_{1}(\mathbb{C}^{n}))$$\pi_{d+e}(W_{r-1}(\mathbb{C}^{n-1}))$$\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$$\pi_{d+e}(W_{1}(\mathbb{C}^{n})).$$\scriptstyle{f_{1}}$$\scriptstyle{f_{2}}$$\scriptstyle{f_{3}}$$\scriptstyle{f_{4}}$
This differs from diagram (9) by the omission of $f_{5}$. Map $f_{1}$ is an isomorphism by Proposition 4.3; the required inequalities may be easily verified with x=n1x=n-1, $y=n$ and $d+1$ playing the part of $d$. Note that inequalities 15 imply the same inequalities with r1r-1, n1n-1 in place of $r$, $n$. Therefore map $f_{2}$ is an injection. Map $f_{4}$ is an isomorphism by Proposition 4.3 again. The result now follows by a diagram chase. ∎
Remark 4.6.
If the Beilinson-Soulé conjecture holds for $k$, then the condition $n-1\leq e$ may be replaced by $\min\{n-1,2n-d\}\leq e$ in each of Theorems 4.4 and 4.5. The proofs are unchanged; the result on which they rely, Proposition 4.3, is stronger.
5. Application to maps of Stiefel varieties and stably free modules
For any $r\in\{1,\dots,n\}$, there is a projection map $\rho:V_{r}(\mathbb{A}^{n})\to V_{1}(\mathbb{A}^{n})$. The problem of determining whether $\rho$ has a right inverse (alternatively termed “a section” since $\rho$ is a fibre bundle) was studied in [9]. Using Theorem 4.5, we can settle all cases of this problem over $\bar{\mathbb{Q}}$.
We recall the definition of the James numbers $b_{q}$, which are also known as the Atiyah–Todd numbers $M_{q}$. Definitions are given in [16], [4], and these definitions are proved to be equal in [1]. Explicitly, they are described by their $p$-adic valuations, $v_{p}$, for all primes $p$:
$v_{p}(b_{q})=\begin{cases}\max\left\{s+v_{p}(s)\>\Big|\>1\leq s\leq\lfloor\frac{q-1}{p-1}\rfloor\right\},\quad\text{ if $q\geq p$;}\\ 0\qquad\text{ otherwise.}\end{cases}$
The first few James numbers may easily be listed:
$b_{2}=2;\quad b_{3}=b_{4}=2^{3}3=24;\quad b_{5}=2^{6}3^{2}5=2880;\quad$
We remark that the James numbers grow quickly. We establish a rough bound: $v_{2}(b_{q})\geq q-1$, so that $b_{q}\geq 2^{q-1}$.
Theorem 5.1.
Let $n$ be a positive integer and suppose $2\leq r\leq n-2$. Then
$\rho:V_{r}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})\to V_{1}(\mathbb{A}_{\bar{\mathbb{Q}}}^{n})$
has a right inverse if and only if $b_{r}\mid n$
Proof.
If $b_{r}\nmid n$, then there cannot be a right inverse by virtue of [20, Thm. 6.5]. Note that this result disallows right inverses even in the $\mathbb{A}^{1}$-homotopy category $\text{\bf H}(\bar{\mathbb{Q}})$.
Let us therefore suppose that $b_{r}\mid n$, and try to construct a right inverse. The case of $r=2$ follows from a well known construction of Bass, first published to our knowledge in [20, Prop. 2.2(b)] as a construction of a free summand in a projective module.
We therefore may suppose $r\geq 3$. By using the rough bound we established earlier, we deduce that $r\leq\log_{2}(n)+1$. If $n\geq 9$, this implies that $r<n/2$. For $n\in\{5,6,7,8\}$, the inequality $b_{r}\leq n$ implies $b_{r}=2$, so $r<n/2$ in this case as well.
Theorem 4.5 applies with (n2,n)(n-2,n) playing the part of $(d,e)$ and (r1,n1)(r-1,n-1) playing the part of $(r,n)$. Inequality 2 of Theorem 4.5 required $r<n/2$ in this case. We deduce that the realization map
$\pi_{n-2+n\alpha}(V_{r-1}(\mathbb{A}^{n-1}))\to\pi_{2n-2}(W_{r-1}(\mathbb{C}^{n-1})$
is injective. We now use [9, Prop. 7.3], which tells us that $\rho$ has a right inverse if the corresponding map of complex Stiefel manifolds
$\rho(\mathbb{C}):W_{r}(\mathbb{C}^{n})\to W_{1}(\mathbb{C}^{n})$
has a right inverse. This has a right inverse by [1, Thm. 1.1]. ∎
Corollary 5.2.
Suppose $R$ is a commutative ring containing an algebraic closure of $\mathbb{Q}$. Let $P$ be an $R$-module for which there exists an isomorphism
$P\oplus R\cong R^{n}$
for some positive integer $n$. If $b_{r}\mid n$, then there exists a decomposition
$P\cong Q\oplus R^{r-1}.$
Proof.
This follows from Theorem 5.1 and [9, Prop. 4.4]. ∎
Appendix A Homological algebra modulo uniquely divisible subgroups
Lemma A.1.
If
$0\to A\to B\to C\to 0$
is a short exact sequence of abelian groups with the property that two of the groups $A,B,C$ are uniquely divisible, then so is the third.
Proof.
By definition, an abelian group $M$ is uniquely divisible if the self-maps $\times n:M\to M$ are isomorphisms for all natural numbers $n$. The result follows easily from the snake lemma. ∎
Proposition A.2.
Consider a commutative diagram of abelian groups
(10) $A_{1}$$A_{2}$$A_{3}$$A_{4}$$A_{5}$$B_{1}$$B_{2}$$B_{3}$$B_{4}$$B_{5},$$\scriptstyle{\phi_{1}}$$\scriptstyle{\phi_{2}}$$\scriptstyle{\phi_{3}}$$\scriptstyle{\phi_{4}}$$\scriptstyle{\phi_{5}}$
in which the rows are exact sequences. Suppose that $\phi_{1},\phi_{2},\phi_{3},\phi_{4}$ are surjective and their kernels are uniquely divisible. Suppose further that $B_{3}$ is torsion. Then $\phi_{3}$ is surjective and $\ker(\phi_{3})$ is uniquely divisible.
Proof.
Treat diagram (10) as a double complex where $A_{i}$ is in degree $(i,0)$ and $B_{i}$ is in degree $(i,1)$. There are two spectral sequences calculating the homology of the total complex: for each of these, the nonzero groups on the $E_{0}$-page coincide with the groups in (10). In one, the $d_{0}$ differentials are the horizontal arrows in (10), and since the rows are exact, we see that the homology of the total complex vanishes in total degrees $3$ and $4$.
The other spectral sequence has as $E_{0}$-page the following diagram
$A_{1}$$A_{2}$$A_{3}$$A_{4}$$A_{5}$$B_{1}$$B_{2}$$B_{3}$$B_{4}$$B_{5},$$\scriptstyle{\phi_{1}}$$\scriptstyle{\phi_{2}}$$\scriptstyle{\phi_{3}}$$\scriptstyle{\phi_{4}}$$\scriptstyle{\phi_{5}}$
so that the $E_{1}$-page is
(11) $\ker(\phi_{1})$$\ker(\phi_{2})$$\ker(\phi_{3})$$\ker(\phi_{4})$$\ker(\phi_{5})$${\operatorname{coker}(\phi_{3})}0{0}0.{0.}$
On the $E_{2}$-page and later, there are no nonzero differentials arriving at or emanating from the $(3,0)$-group, which must be 0 by the $E_{\infty}$-page, so that we deduce that the upper sequence in (11) is exact at $\ker(\phi_{3})$. The terms in this sequence, other than $\ker(\phi_{3})$, are known to be uniquely divisible, so that the $5$-lemma implies that $\ker(\phi_{3})$ is also uniquely divisible.
The $E_{2}$-page takes the form
[…]
where $Q$ is the cokernel of $\ker(\phi_{4})\to\ker(\phi_{5})$ and $\ast$ denotes some unspecified abelian groups. For dimensional reasons, there are no further nonzero differentials, from which it follows that $\operatorname{coker}(\phi_{3})\to Q$ must be an injection: the abutment of the sequence is 0 in total degree $4$.
Since $\ker(\phi_{4})$ is uniquely divisible, its image in $\ker(\phi_{5})$ is divisible, and therefore uniquely divisible. It follows from Lemma A.1 that $Q$ is uniquely divisible. In particular, the map $\operatorname{coker}(\phi_{3})\to Q$, which has a torsion domain, must be 0. Since it is also an injection, we deduce that $\operatorname{coker}(\phi_{3})$ is 0 as required. ∎